# TransVIP: Speech to Speech Translation System with Voice and Isochrony Preservation

### v2

@@ 1 Introduction
To address these issues, we propose TransVIP, a speech-to-speech Translation framework with Voice and Isochrony Preservation. TransVIP employs a consecutive generation approach, simplifying the complex S2ST task into two sequential tasks while maintaining an end-to-end framework. The generation of the target speech is conditioned not only on the semantic information, as in conventional S2ST models, but also on the isochrony and acoustic information derived from the source speech. The corresponding overview of TransVIP is demonstrated in Figure 1. We evaluate the performance of TransVIP for French-English mutual translation using a subset of the CVSS-T test set [8]. The results demonstrate that TransVIP outperforms the publicly available SOTA models such as a larger Seamless Expressive model. The generated audio and video samples are available at https://aka.ms/transvip. The main contributions of the paper can be summarized as follows:
@@
    We propose to disentangle various information required to learn in the training stage through employing separated encoders. It can enhance the transfer of voice characteristics and isochrony temporal alignment from the source to the target speech in the translation process. Additionally, it facilitates the design of lightweight modules for more effective individual information learning.
@@ 3.4 Multi-task Training
We claim that this translation model is a cascaded trainable end-to-end model. Because the distribution $P(Y^{\prime}|S,A,I,T)$ and $P(T|S,I)$ can be fitted both separately and jointly using various dataset. It significantly enhances data accessibility, which is crucial given the scarcity of S2ST datasets. We utilize these datasets during training as follows:
@@ 4.1 Nerual Codec with Semantic Distillation
Our neural codec is designed to provide high reconstruction quality and compatibility with language model predictions. We refer to our codec as the multi-lingual Semantic-Aware Speech Codec (SASC), which integrates two existing methodologies: DAC [22] and SpeechTokenizer. DAC employs innovative factorized and L2-normalized codes in vector quantization and finely tuned hyperparameters to achieve state-of-the-art reconstruction quality. SpeechTokenizer, on the other hand, uses a Hubert encoder as a teacher and performs semantic distillation between it and the first-layer Residual Vector Quantization (RVQ) codes. This process enriches the RVQ codes with additional semantic information, facilitating downstream language modeling tasks.
@@ 4.2 Textless Non-autoregressive Acoustic Modeling
Our approach diverges from prior studies by excluding textual or phonetic inputs, rendering it a completely textless model. Previous research indicated that omitting semantic information such as text or phonemes typically significantly decreases intelligibility. However, our experiments demonstrate that the textless model either matches or surpasses the performance of models with text inputs, owing to semantic distillation within the SASC. A comparison of text and textless model is shown in appendix 6.3
@@ 5.1 Experimental Settings
We use at most, a 10-second prompt for the joint translation model and a 5-second prompt for the NAR acoustic model. If the source speech is shorter than that limitation we just use the whole utterance as the prompt. For baseline, we compared our model with a textless S2UT model and two Seamless models. The textless S2UT model is the All-to-English XM Transformer [^6] https://github.com/facebookresearch/fairseq/tree/ust/examples/speech_matrix release by SpeechMatrix[49]. For seamless models, We compared with two versions, the medium version, and the expressive version. The medium version uses the same speech-to-text translation model as the one we use to initialize our joint translation model, while the expressive version has larger model size and a stronger text-to-speech model. For codec evaluation, we employ LibriSpeech [50] test-clean, which is widely adopted for codec and TTS evaluation. We compare our SASC with Encodec [21], Speechtokenizer [9] and DAC[22]. We evaluate S2ST performance on translation(BLEU), speaker&prosody similarity(SIM & AutoPCP), Isochrony control(Rate, Pause & SLC), and naturalness.
@@ 5.2 S2ST Evaluation Result
We observed that the quality of the conditioning speech prompt significantly influences the naturalness of the generated speech, as measured by the NISQA score. For instance, when processing low-quality real French data (Fr-En), SeamlessExpressive registers the lowest score at 2.91. Conversely, with high-quality synthesized English data (En-Fr) as input, SeamlessExpressive achieves the highest score of 3.57. Meanwhile, our TransVIP model demonstrates consistent performance, scoring 3.19 in the Fr-En direction and 3.40 in the en-fr direction. Given that the NISQA score can present difficulties in evaluating the quality or naturalness of noisy speech, we also assess the naturalness of speech using a subjective metric known as the Comparative Mean Opinion Score (CMOS). For this evaluation, 12 professional linguistic experts judged 120 pairs of utterances translated by TransVIP and SeamlessExpressive. The results indicated that TransVIP’s performance is significantly better (p<0.01) than SeamlessExpressive, i.e., TransVIP outperforms SeamlessExpressive by achieving a preference score higher by 0.46. The instruction for the subjective test is shown in Appendix G.
@@ 5.3 Codec Evaluation Result
We show the evaluation result in Table 6 and group the result by bandwidth(BW). The SpeechTokenizer that has undergone semantic distillation performs well in the NISQA score, producing a clean and natural voice. On the other hand, DAC performs well in similarity and WER, showing a strong capability of a faithful reconstruction of the original speech. Finally, our SASC successfully combines the strength of two codecs, showing a very competitive result in all four metrics.
@@ 6.3 Ablation Study on the NAR Acoustic Model
Table 5: Ablation study on the NAR Acoustic model and its inference strategy. The BLEU score is speech BLEU(ASR-BLEU). Abbreviation: LL(Librilight), LH(Libriheavy), SA(SeamlessAlign), CV(Common Voice 15), VP(VoxPopuli)
NAR Model Datasets SIM ASR-BLEU Nat.
NAR w/o text LL,VP,SA,CV 0.320 32.60 3.19
@@ 7 Conclusion, Limitations and future works
Broader Impact: TransVIP was developed to improve cross-lingual communication and assist with automated video dubbing workflows. However, it carries potential risks, including the misuse of the model for impersonating specific speakers by inputting targeted text and speaker prompts. Additionally, as a generative model, there is a possibility it may produce toxic or biased outputs, although such occurrences were not observed in our test samples. It is imperative to monitor for any misuse or harmful outputs rigorously. We must also allow users to report such incidents, ensuring responsible usage and continuous system improvement.
@@
We extend our gratitude to Chung-Hsien Tsai, Canrun Li, and Zhen Xiao for their invaluable contribution in conducting the subjective evaluation and for their efforts in assembling the video dubbing material. Also, we appreciate Microsoft Azure for providing computing services and GPUs.
@@ Appendix B Implementaion Details
The primary encoder-decoder (Enc-Dec) architecture of the joint model is initialized using the SeamlessM4T S2T model, while the other submodules are trained from scratch. The first-layer RVQ codes are integrated into the vocabulary of the foundational S2T model. As mentioned in Section 3.4, this model is trained using multiple datasets, including two S2ST datasets: CVSS-T[44] and SeamlessAlign [^11] https://github.com/facebookresearch/seamless_communication/blob/main/docs/m4t/seamless_align_README.md, one internal ST dataset, and one ASR dataset: Common Voice[45] version 15 (English and French subsets). An offline preprocessing step is performed to convert the speech labels into a discrete codec form. During this process, we utilize a VAD tool [^12] https://github.com/snakers4/silero-vad, MIT License to eliminate silence at the beginning and end of the speech recordings. This VAD tool is also used to generate the isochrony embeddings during both the training and inference phases. The input to the speech semantic encoder is Fbank features and text is tokenized into BPE, which is consistent with the Seamless framework. The acoustic encoder is a six-layer standard Transformer encoder with a hidden size of 1024. Its input features are extracted from raw waveforms by the SASC encoder, providing richer acoustic information than Fbank features. For decoding we employed a beam search algorithm with a beam size set to 5.
@@
SASC Model
We train the model on audio with a frequency of 16k Hz. The downsample rate is set at 320, resulting in 50 tokens/frames per second, instead of the 75 tokens used in the official DAC checkpoint. This adjustment is primarily made to ensure a consistent token rate that matches the semantic token rate, thereby facilitating the distillation process. The model is trained with 24 RVQ codebooks while only 16 are employed in the translation system. The SASC model is trained using the full set of the Common Voice version 15 corpus, which includes hours of multilingual speech data. During each training iteration, a random 2-second segment is cropped from the speech samples.
@@ Appendix C Additional Ablation on the Choice of Codec
In this section, we compare training TransVIP using different codecs: SpeechTokenizer [9] and DAC [22]. In this ablation study, the joint translation model is trained with a subset containing only CVSS-T Fr-En uni-direction data. For the NAR acoustic model, both DAC and SASC use 16 codec layers, while SpeechTokenizer uses 8 layers, as it only has an 8-layer version. The results are shown in Table 6. Compared to SpeechTokenizer, the model trained with SASC exhibits superior performance in all aspects. Most notably, the speaker similarity improved by 0.04, from 0.226 to 0.264, aligning with the improvement in codec re-synthesis results.
@@
Additionally, we find that when trained with DAC, the joint translation model fails to generate reasonable output even when the training loss appears to converge. To investigate this, we plot the validation loss curve of the joint translation model in Figure 3. It shows that SpeechTokenizer has the lowest loss, followed by SASC, and then DAC. Furthermore, we observed that SpeechTokenizer can generate reasonable output with minimal training steps. We attribute this finding to two reasons: 1) SpeechTokenizer contains most semantic information in its first-layer codec. Although SASC also undergoes semantic distillation, it still contains acoustic information in the first layer, trading for better speech reconstruction quality. 2) The enhancement of RVQ in DAC, which is also utilized in SASC, leads to an even utilization rate across the codebook. This could pose a challenge for language modeling. However, our SASC method effectively balances the quality of reconstruction and the affinity of the language model.
@@ Appendix G Subjective Evaluation
The instruction of CMOS naturalness test is shown in Figure 4. Much better/better/slightly better/can’t tell equals to 3/2/1/0 score respectively.

### v3

@@ 1 Introduction
To address these issues, we propose TransVIP, a speech-to-speech Translation framework with Voice and Isochrony Preservation. TransVIP employs a consecutive generation approach, simplifying the complex S2ST task into two sequential tasks while maintaining an end-to-end framework. The generation of the target speech is conditioned not only on the semantic information, as in conventional S2ST models, but also on the isochrony and acoustic information derived from the source speech. The corresponding overview of TransVIP is demonstrated in Figure 1. We evaluate the performance of TransVIP for French-English mutual translation using a subset of the CVSS-T test set [8]. The results demonstrate that TransVIP outperforms the publicly available SOTA models such as a larger Seamless Expressive model. The generated audio samples are available at https://aka.ms/transvip. The training code and script are available at https://github.com/nethermanpro/transvip. The main contributions of the paper can be summarized as follows:
@@
    We propose to disentangle various information required to learn in the training stage by employing separated encoders. It can enhance the transfer of voice characteristics and isochrony temporal alignment from the source to the target speech in the translation process. Additionally, it facilitates the design of lightweight modules for more effective individual information learning.
@@ 3.4 Multi-task Training
We claim that this translation model is a cascaded trainable end-to-end model. Because the distribution $P(Y^{\prime}|S,A,I,T)$ and $P(T|S,I)$ can be fitted both separately and jointly using various datasets. It significantly enhances data accessibility, which is crucial given the scarcity of S2ST datasets. We utilize these datasets during training as follows:
@@ 4.1 Nerual Codec with Semantic Distillation
Our neural codec is designed to provide high reconstruction quality and compatibility with language model predictions. We refer to our codec as the Semantic-Aware Speech Codec (SASCodec), which integrates two existing methodologies: DAC [22] and SpeechTokenizer. DAC employs innovative factorized and L2-normalized codes in vector quantization and finely tuned hyperparameters to achieve state-of-the-art reconstruction quality. SpeechTokenizer, on the other hand, uses a Hubert encoder as a teacher and performs semantic distillation between it and the first-layer Residual Vector Quantization (RVQ) codes. This process enriches the RVQ codes with additional semantic information, facilitating downstream language modeling tasks.
@@ 4.2 Textless Non-autoregressive Acoustic Modeling
Our approach diverges from prior studies by excluding textual or phonetic inputs, rendering it a completely textless model. Previous research indicated that omitting semantic information such as text or phonemes typically significantly decreases intelligibility. However, our experiments demonstrate that the textless model either matches or surpasses the performance of models with text inputs, owing to semantic distillation within the SASCodec.
@@ 5.1 Experimental Settings
We use at most, a 10-second prompt for the joint translation model and a 5-second prompt for the NAR acoustic model. If the source speech is shorter than that limitation we just use the whole utterance as the prompt. For baseline, we compared our model with a textless S2UT model and two Seamless models. The textless S2UT model is the All-to-English XM Transformer [^6] https://github.com/facebookresearch/fairseq/tree/ust/examples/speech_matrix release by SpeechMatrix[49]. For seamless models, We compared with two versions, the medium version, and the expressive version. The medium version uses the same speech-to-text translation model as the one we use to initialize our joint translation model, while the expressive version has larger model size and a stronger text-to-speech model. For codec evaluation, we employ LibriSpeech [50] test-clean, which is widely adopted for codec and TTS evaluation. We compare our SASCodec with Encodec [21], Speechtokenizer [9] and DAC[22]. We evaluate S2ST performance on translation(BLEU), speaker&prosody similarity(SIM & AutoPCP), Isochrony control(Rate, Pause & SLC), and naturalness. Please refer to Appendix A for the details of evaluation matrics.
@@ 5.2 S2ST Evaluation Result
We observed that the quality of the conditioning speech prompt significantly influences the naturalness of the generated speech, as measured by the NISQA score. For instance, when processing low-quality real French data (Fr-En), SeamlessExpressive registers the lowest score at 2.91. Conversely, with high-quality synthesized English data (En-Fr) as input, SeamlessExpressive achieves the highest score of 3.57. Meanwhile, our TransVIP model demonstrates consistent performance, scoring 3.19 in the Fr-En direction and 3.40 in the en-fr direction.
@@ 5.3 Codec Evaluation Result
We show the evaluation result in Table 2 and group the result by bandwidth(BW). The SpeechTokenizer that has undergone semantic distillation performs well in the NISQA score, producing a clean and natural voice. On the other hand, DAC performs well in similarity and WER, showing a strong capability of a faithful reconstruction of the original speech. Finally, our SASCodec successfully combines the strength of two codecs, showing a very competitive result in all four metrics.
@@ 6.3 Ablation Study on the NAR Acoustic Model
Table 5: Ablation study on the NAR Acoustic model and its inference strategy.
NAR Model SIM ASR-BLEU Nat.
NAR w/o text 0.320 32.60 3.19
@@ 7 Conclusion, Limitations and future works
Broader Impact: TransVIP was developed to improve cross-lingual communication and assist with automated video dubbing workflows. However, it carries potential risks, including the misuse of the model for impersonating specific speakers by inputting targeted text and speaker prompts. Additionally, as a generative model, there is a possibility it may produce toxic or biased outputs, although such occurrences were not observed in our test samples.
@@
We extend our gratitude to Chung-Hsien Tsai, Canrun Li, and Zhen Xiao for their invaluable contribution in conducting the subjective evaluation and for their efforts in assembling the video dubbing material.
@@ Appendix B Implementaion Details
The primary encoder-decoder (Enc-Dec) architecture of the joint model is initialized using the SeamlessM4T S2T model, while the other submodules are trained from scratch. The first-layer RVQ codes are integrated into the vocabulary of the foundational S2T model. As mentioned in Section 3.4, this model is trained using multiple datasets, including two S2ST datasets: CVSS-T[44] and SeamlessAlign [^11] https://github.com/facebookresearch/seamless_communication/blob/main/docs/m4t/seamless_align_README.md, one internal ST dataset, and one ASR dataset: Common Voice[45] version 15 (English and French subsets). An offline preprocessing step is performed to convert the speech labels into a discrete codec form. During this process, we utilize a VAD tool [^12] https://github.com/snakers4/silero-vad, MIT License to eliminate silence at the beginning and end of the speech recordings. This VAD tool is also used to generate the isochrony embeddings during both the training and inference phases. The input to the speech semantic encoder is Fbank features and text is tokenized into BPE, which is consistent with the Seamless framework. The acoustic encoder is a six-layer standard Transformer encoder with a hidden size of 1024. Its input features are extracted from raw waveforms by the SASCodec encoder, providing richer acoustic information than Fbank features. For decoding we employed a beam search algorithm with a beam size set to 5.
@@
SASCodec Model
We train the model on audio with a frequency of 16k Hz. The downsample rate is set at 320, resulting in 50 tokens/frames per second, instead of the 75 tokens used in the official DAC checkpoint. This adjustment is primarily made to ensure a consistent token rate that matches the semantic token rate, thereby facilitating the distillation process. The model is trained with 24 RVQ codebooks while only 16 are employed in the translation system. The SASCodec model is trained using the full set of the Common Voice version 15 corpus, which includes hours of multilingual speech data. During each training iteration, a random 2-second segment is cropped from the speech samples.
@@ Appendix C Additional Ablation on the Choice of Codec
In this section, we compare training TransVIP using different codecs: SpeechTokenizer [9] and DAC [22]. In this ablation study, the joint translation model is trained with a subset containing only CVSS-T Fr-En uni-direction data. For the NAR acoustic model, both DAC and SASCodec use 16 codec layers, while SpeechTokenizer uses 8 layers, as it only has an 8-layer version. The results are shown in Table 6. Compared to SpeechTokenizer, the model trained with SASCodec exhibits superior performance in all aspects. Most notably, the speaker similarity improved by 0.04, from 0.226 to 0.264, aligning with the improvement in codec re-synthesis results.
@@
Additionally, we find that when trained with DAC, the joint translation model fails to generate reasonable output even when the training loss appears to converge. To investigate this, we plot the validation loss curve of the joint translation model in Figure 3. It shows that SpeechTokenizer has the lowest loss, followed by SASCodec, and then DAC. Furthermore, we observed that SpeechTokenizer can generate reasonable output with minimal training steps. We attribute this finding to two reasons: 1) SpeechTokenizer contains most semantic information in its first-layer codec. Although SASCodec also undergoes semantic distillation, it still contains acoustic information in the first layer, trading for better speech reconstruction quality. 2) The enhancement of RVQ in DAC, which is also utilized in SASCodec, leads to an even utilization rate across the codebook. This could pose a challenge for language modeling. However, our SASCodec method effectively balances the quality of reconstruction and the affinity of the language model.
@@ Appendix G Subjective Evaluation
Given that the NISQA score can present difficulties in evaluating the quality or naturalness of noisy speech, we also assess the naturalness of speech using a subjective metric known as the Comparative Mean Opinion Score (CMOS). For this evaluation, 12 professional linguistic experts judged 120 pairs of utterances translated by TransVIP and SeamlessExpressive. The results indicated that TransVIP’s performance is significantly better (p<0.01) than SeamlessExpressive, i.e., TransVIP outperforms SeamlessExpressive by achieving a preference score higher by 0.46. The instruction of CMOS naturalness test is shown in Figure 4. Much better/better/slightly better/can’t tell equals to 3/2/1/0 score respectively.